OpenStudy (anonymous):

medal & fan :) Which set of line segments could create a right triangle? 15, 30, 35 15, 36, 39 15, 20, 29 5, 15, 30

OpenStudy (anonymous):

Pythagorean theorem.

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

First choice. What is 15^2 and 30^2?

OpenStudy (horsegirl27):

I think A

OpenStudy (anonymous):

No direct answers.

OpenStudy (some.random.cool.kid):

plus i dont think its A

OpenStudy (anonymous):

when you guys are done here, mind helping me? I need someone to check my answers

OpenStudy (horsegirl27):

Oops sorry:)

OpenStudy (anonymous):

Thank you

OpenStudy (horsegirl27):

I'm not sure if it is A but I think it might be

OpenStudy (theedwardsfamily):

get a calculator and evaluate the following: 15^2 + 36^2 = 39^2